OpenStudy (anonymous):

I am 56 trying to go back to school can someone tell me what this is: 82+9/3-5=

OpenStudy (anonymous):

80

OpenStudy (anonymous):

begin with the division first: 82+(9/3)-5=(82+3)-5=85-5=80 remember the MDAS rule (performing multiplication or division first before addition and subtraction) it will help if you use grouping symbols such as parentheses to avoid confusion. and learn associativity and distributivity properties of the operations :)

OpenStudy (anonymous):

PEMDAS- P- get rid of parenthesis E- get rid of exponets M- multiplication D- Division A- Addtion S-Subtraction.... Follow these rules and you will be okay,,,best of luck to you!! I too have gone back to school after a very long time